Colophon: About HtX

The modules contained in this book were originally prepared for distribution via the World Wide Web using the HtX development system.

HtX uses a L^{A}TEX-like source file to specify a set of interlinked HTML pages. These are formatted according to a style specified by the author. Authors can create their own styles or extend and/or overwrite existing styles to produce intricate documents with minimal effort. HtX’s meta-language provides flexibility and extensibility in expressing style apart from content.

HtX’s power lies in its ability to provide the pieces for creating a uniform style. Rather than enforcing any particular visual look and feel, or pattern of interconnection, HtX allows the author to build his or her own style using the primitive tools provided by the system.

An HtX source file can be easily converted to printable form, due to the similarity of HtX and L^{A}TEX. This conversion procedure was employed in the preparation of this book.

HtX 1.0 styles exist for creating laboratory manuals, lectures and lecture notes, class syllabi, and general purpose documents. The updated styles of HtX 1.1 make extensive use of newer browser technology, such as frames. A WYSIWYG interface for HtX, called WebFaces, is currently under development.

For more information, including downloading instructions, see the HtX home page:

                     http://www.cs.oberlin.edu/htx/